### v1

# Property $\mathrm{(TTT)}$ and the Cowling–Haagerup constant

Abstract.
We show that Property $\mathrm{(TTT)}$ is an obstruction to weak amenability with Cowling–Haagerup constant $1$. More precisely, if $G$ is a countable group and $H$ is an infinite subgroup of $G$ such that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$, then the weak Haagerup constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ is strictly greater than $1$. We apply this result to some semidirect products and lattices in higher rank algebraic groups.
Key words and phrases:
Property (TTT), weak amenability, weak Haagerup property
2020 Mathematics Subject Classification:
Primary 22D55; Secondary 43A07, 46L07
This work is supported by the FONDECYT project 3230024.
1. Introduction
A countable group $G$ is said to be amenable if there is a sequence of finitely supported, positive definite functions $\varphi_{n}:G\to\mathbb{C}$ such that
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G.$
More generally, we say that $G$ is weakly amenable if there is a sequence of finitely supported functions $\varphi_{n}:G\to\mathbb{C}$ and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G,$
and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C,$
where $B_{2}(G)$ stands for the space of Herz–Schur multipliers of $G$; see Section 2.2. The Cowling–Haagerup constant $\boldsymbol{\Lambda}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds. This is indeed a weakening of amenability because every positive definite function $\varphi$ satisfies $\|\varphi\|_{B_{2}(G)}=\varphi(1)$. Hence, if $G$ is amenable, then $\boldsymbol{\Lambda}(G)=1$. If $G$ is not weakly amenable, we simply set $\boldsymbol{\Lambda}(G)=\infty$.
A strong negation of amenability is given by Property $\mathrm{(T)}$. Let $G$ be a countable group, and let $H$ be a subgroup of $G$. We say that the pair $(G,H)$ has relative Property $\mathrm{(T)}$ if every sequence of positive definite functions on $G$ converging pointwise to $1$ converges uniformly on $H$. If this condition holds for $H=G$, we simply say that $G$ has Property $\mathrm{(T)}$. We refer the reader to [1] for a thorough account on Property $\mathrm{(T)}$ and its different characterisations. It follows from the definitions that, if $H$ is infinite and the pair $(G,H)$ has relative Property $\mathrm{(T)}$, then $G$ cannot be amenable.
For weak amenability, however, the situation is quite different, and Property $\mathrm{(T)}$ is no longer an obstruction. This was first observed by Cowling and Haagerup [2], who showed that the rank 1 Lie groups $\operatorname{Sp}(n,1)$ and $\operatorname{F}_{4,-20}$, as well as their lattices, are weakly amenable. Afterwards, the list of examples was considerably expanded by Ozawa [8], who showed that all hyperbolic groups are weakly amenable. It must be noted that, in a sense that can be made very precise, “almost all” hyperbolic groups have Property $\mathrm{(T)}$; see [11] and [7] for details. In spite of these facts, the following question remains open.
Question 1.1.
Does there exist a pair $(G,H)$ with relative Property $\mathrm{(T)}$ such that $H$ is infinite and $\boldsymbol{\Lambda}(G)=1$?
As a matter of fact, all the known examples of groups with $\boldsymbol{\Lambda}(G)=1$ satisfy the Haagerup property, meaning that there is a sequence of positive definite functions $\varphi_{n}:G\to\mathbb{C}$ such that $\varphi_{n}$ vanishes at infinity for all $n$, and
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G.$
The problem of determining whether this implication holds in general is often referred to as Cowling’s conjecture. In order to attack this problem, Knudby [5] introduced the weak Haagerup property as a generalisation of both weak amenability and the Haagerup property. In this case, the approximation of the identity is given by functions vanishing at infinity and which are uniformly bounded in $B_{2}(G)$. We can also associate a constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ to this property, which is defined in similar fashion; see Section 2.2. We always have
$\boldsymbol{\Lambda}_{\mathrm{WH}}(G)\leq\boldsymbol{\Lambda}(G).$
In this paper, we address a variation of Question 1.1 by considering a stronger version of Property $\mathrm{(T)}$. In [9], Ozawa introduced Property $\mathrm{(T_{P})}$ as a non-equivariant version of Property $\mathrm{(T)}$, where positive definite functions are replaced by positive definite kernels, with some additional assumptions; see Section 2.3 for details. An a priori weaker property, called Property $\mathrm{(TTT)}$, which was also introduced in [9], is defined in terms of boundedness of wq-cocycles; see Section 2.3. Very recently, Dumas [3] showed that Property $\mathrm{(T_{P})}$ and Property $\mathrm{(TTT)}$ are equivalent. Using this fact, we prove the following.
Theorem 1.2.
Let $G$ be a countable group, and let $H$ be an infinite subgroup of $G$ such that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$. Then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ belongs to $(1,\infty]$.
In [3, Corollary 4.8], Property $\mathrm{(TTT)}$ was established for lattices in higher rank algebraic groups. As a consequence, we obtain the following. We refer the reader to [4] for details on algebraic groups.
Corollary 1.3.
Let $\mathbb{K}$ be a local field, and let $G$ be a connected, almost $\mathbb{K}$-simple $\mathbb{K}$-group with $\operatorname{rank}_{\mathbb{K}}G\geq 2$. If $\Gamma$ is a lattice in $G(\mathbb{K})$, then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ belongs to $(1,\infty]$.
Property $\mathrm{(TTT)}$ is strictly stronger than Property $\mathrm{(T)}$, as witnessed by the fact that hyperbolic groups are a-$\mathrm{TTT}$-menable, which is a strong negation of Property $\mathrm{(TTT)}$; see [9]. However, these properties are equivalent for some semidirect products; see [9, Proposition 3]. Using this fact, we obtain the following.
Corollary 1.4.
Let $G$ be a countable group such that $G=\Gamma\ltimes A$, where $A$ is infinite and abelian. If the pair $(G,A)$ has relative Property $\mathrm{(T)}$, then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ belongs to $(1,\infty]$.
2. Main section
2.1. Schur multipliers
We begin by reviewing some facts about Schur multipliers and positive definite kernels; for more detailed presentations, we refer the reader to [10, Chapter 5] and [1, Appendix C]. Let $X$ be a set. We say that $\theta:X\times X\to\mathbb{C}$ is a Schur multiplier on $X$ if there is a Hilbert space $\mathcal{H}$ and bounded maps $P,Q:X\to\mathcal{H}$ such that
$\theta(x,y)=\langle P(x),Q(y)\rangle,\quad\forall x,y\in X.$ (1)
We denote by $V(X)$ the space of Schur multipliers on $X$, and we endow it with the norm
$\|\theta\|_{V(X)}=\inf\left(\sup_{x\in X}\|P(x)\|\right)\left(\sup_{y\in X}\|Q(y)\|\right),$
where the infimum is taken over all the decompositions as in (1). In the particular case in which (1) can be obtained with $P=Q$, we say that $\theta$ is a positive definite kernel on $X$. In addition, we say that $\theta$ is normalised if $\theta(x,x)=1$ for all $x\in X$. We will make use of the following basic fact; see [1, Theorem C.3.2].
Proposition 2.1.
Let $\xi:X\to\mathcal{H}$ be a map from a set $X$ to a Hilbert space $\mathcal{H}$. For every $t>0$, the kernel
$(x,y)\in X\times X\ \longmapsto\ e^{-t\|\xi(x)-\xi(y)\|^{2}}$
is positive definite.
2.2. Weak amenability and the weak Haagerup property
Now let $G$ be a group. We say that $\varphi:G\to\mathbb{C}$ is a Herz–Schur multiplier if the map
$(x,y)\in G\times G\ \longmapsto\ \varphi(y^{-1}x)$
is a Schur multiplier on $G$. We denote by $B_{2}(G)$ the space of Herz–Schur multipliers on $G$, and we endow it with the norm inherited from $V(G)$, for which it becomes a Banach algebra. Recall that $G$ is weakly amenable if there is a sequence of finitely supported functions $\varphi_{n}:G\to\mathbb{C}$ and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\varphi_{n}(x)=1,\quad\forall x\in G,$
and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C.$
The Cowling–Haagerup constant $\boldsymbol{\Lambda}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds.
More generally, $G$ has the weak Haagerup property if there is a sequence of functions $\varphi_{n}:G\to\mathbb{C}$ vanishing at infinity, and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\varphi_{n}(x)=1,\quad\forall x\in G,$
and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C.$
The weak Haagerup constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds. It readily follows that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)\leq\boldsymbol{\Lambda}(G)$. We point out that these constants may very well be different. Even more, there are groups $G$ such that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$ and $\boldsymbol{\Lambda}(G)=\infty$. One example of such a group is the wreath product $\mathbb{Z}\wr\mathbb{F}_{2}$, where $\mathbb{F}_{2}$ is the free group on 2 generators; see [6] for details.
The weak Haagerup property was introduced in [5] with the goal of studying the connection between weak amenability and the Haagerup property. The following result will be essential to our purposes; see [5, Theorem 1.2] and [5, Proposition 4.3].
Theorem 2.2 (Knudby).
Let $G$ be a countable group with $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$. Then there exist a function $\psi:G\to[0,\infty)$, a Hilbert space $\mathcal{H}$, and maps $R,S:G\to\mathcal{H}$ such that
$\psi(y^{-1}x)=\|R(x)-R(y)\|^{2}+\|S(x)+S(y)\|^{2},\quad\forall x,y\in G,$ (2)
and $\lim_{x\to\infty}\psi(x)=\infty$.
Observe that the map $S:G\to\mathcal{H}$ given by Theorem 2.2 takes values in the sphere of radius $\frac{1}{2}\sqrt{\psi(1)}$.
2.3. Property $\mathrm{(T_{P})}$ and Property $\mathrm{(TTT)}$
Now we discuss two strengthenings of Property $\mathrm{(T)}$ introduced in [9]. Let $G$ be a group and let $\theta:G\times G\to\mathbb{C}$. For every $s\in G$, we will denote by $s\cdot\theta$ the map
$s\cdot\theta(x,y)=\theta(s^{-1}x,s^{-1}y),\quad\forall x,y\in G.$
Now let $H$ be a subgroup of $G$. We say that the pair $(G,H)$ has relative Property $\mathrm{(T_{P})}$ if, for every $\varepsilon>0$, there is a finite subset $K\subset G$ and $\delta>0$ such that, for every normalised positive definite kernel $\theta:G\times G\to\mathbb{C}$ satisfying
$\sup_{s\in G}\|s\cdot\theta-\theta\|_{V(G)}$ $<\delta$ and $\sup_{x^{-1}y\in K}|\theta(x,y)-1|<\delta,$
one has
$\sup_{x,y\in H}|\theta(x,y)-1|<\varepsilon.$
Let $\mathcal{H}$ be a Hilbert space and let $\mathbf{U}(\mathcal{H})$ denote its unitary group. We say that $b:G\to\mathcal{H}$ is a wq-cocycle if there is a map $\pi:G\to\mathbf{U}(\mathcal{H})$ (not necessarily a representation) such that
$\sup_{x,y\in G}\|b(xy)-\pi(x)b(y)-b(x)\|<\infty.$
We say that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$ if every wq-cocycle on $G$ is bounded on $H$. If $H=G$, we simply say that $G$ has Property $\mathrm{(TTT)}$. The following was proved in [9, Theorem 1] and [3, Theorem 2.7].
Theorem 2.3 (Ozawa, Dumas).
Let $G$ be a countable group, and let $H$ be a subgroup of $G$. Then the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$ if and only if it has relative Property $\mathrm{(T_{P})}$.
2.4. Proofs
Now we prove Theorem 1.2. The proof is inspired by that of [9, Theorem 1].
Proof of Theorem 1.2.
Let $G$ be a countable group, and let $H$ be an infinite subgroup of $G$ such that $(G,H)$ has relative Property $\mathrm{(TTT)}$. By Theorem 2.3, $(G,H)$ has relative Property $\mathrm{(T_{P})}$. Assume by contradiction that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$, and let $R,S:G\to\mathcal{H}$ be the maps given by Theorem 2.2. For every $n\geq 1$, we define $\theta_{n}:G\times G\to(0,1]$ by
$\theta_{n}(x,y)=e^{-\frac{1}{n}\|R(x)-R(y)\|^{2}},\quad\forall x,y\in G.$
By Proposition 2.1, $\theta_{n}$ is a normalised positive definite kernel for every $n\geq 1$. Observe that we can view $\mathcal{H}$ as a real Hilbert space with inner product $\mathrm{Re}(\langle\cdot,\cdot\rangle)$. Now consider the full Fock Hilbert space
$\mathcal{F}=\bigoplus_{n\geq 0}\mathcal{H}^{\otimes n},$
where $\mathcal{H}^{\otimes 0}={\mathbb{R}}$ and $\mathcal{H}^{\otimes(n+1)}=\mathcal{H}\otimes\mathcal{H}^{\otimes n}$. The exponential map $\operatorname{EXP}:\mathcal{H}\to\mathcal{F}$ is given by
$\operatorname{EXP}(\xi)=1\oplus\frac{\xi}{\sqrt{1!}}\oplus\frac{\xi\otimes\xi}{\sqrt{2!}}\oplus\frac{\xi\otimes\xi\otimes\xi}{\sqrt{3!}}\oplus\cdots,\quad\forall\xi\in\mathcal{H}.$
We can define a map from $\mathcal{H}$ to the unit sphere of $\mathcal{F}$ by
$\operatorname{E}(\xi)=e^{-\|\xi\|^{2}}\operatorname{EXP}(\sqrt{2}\xi),\quad\forall\xi\in\mathcal{H}.$
This map satisfies
$\langle\operatorname{E}(\xi),\operatorname{E}(\eta)\rangle=e^{-\|\xi-\eta\|^{2}},\quad\forall\xi,\eta\in\mathcal{H}.$
For each $n\geq 1$, define $P_{n},Q_{n}:G\to\mathcal{F}$ by
$P_{n}(x)$ $=\operatorname{E}\left(\frac{1}{\sqrt{n}}S(x)\right),$ $Q_{n}(x)$ $=\operatorname{E}\left(-\frac{1}{\sqrt{n}}S(x)\right),$
for all $x\in G$. Recall the identity (2) and observe that, for all $s,x,y\in G$, $n\geq 1$,
$\theta_{n}(x,y)-s\cdot\theta_{n}(x,y)$ $=e^{-\frac{1}{n}\|R(x)-R(y)\|^{2}}-e^{-\frac{1}{n}\|R(s^{-1}x)-R(s^{-1}y)\|^{2}}$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\left(e^{\frac{1}{n}\|S(x)+S(y)\|^{2}}-e^{\frac{1}{n}\|S(s^{-1}x)+S(s^{-1}y)\|^{2}}\right)$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\left(\langle P_{n}(x),Q_{n}(y)\rangle-\langle P_{n}(s^{-1}x),Q_{n}(s^{-1}y)\rangle\right)$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\langle P_{n}(x)-P_{n}(s^{-1}x),Q_{n}(y)\rangle$
$\quad+e^{-\frac{1}{n}\psi(y^{-1}x)}\langle P_{n}(s^{-1}x),Q_{n}(y)-Q_{n}(s^{-1}y)\rangle.$
Furthermore,
$\|P_{n}(x)-P_{n}(s^{-1}x)\|^{2}$ $=2-2\langle P_{n}(x),P_{n}(s^{-1}x)\rangle$
$=2-2e^{\frac{1}{n}\|S(x)-S(s^{-1}x)\|^{2}}$
$\leq 2-2e^{\frac{1}{n}\psi(1)}.$
Similarly, we get
$\|Q_{n}(y)-Q_{n}(s^{-1}y)\|^{2}\leq 2-2e^{\frac{1}{n}\psi(1)}.$
On the other hand, by [5, Proposition 4.3],
$\left\|(x,y)\mapsto e^{-\frac{1}{n}\psi(y^{-1}x)}\right\|_{V(G)}\leq 1.$
All this shows that
$\|s\cdot\theta_{n}-\theta_{n}\|_{V(G)}\leq 2\sqrt{2}\left(1-e^{\frac{1}{n}\psi(1)}\right)^{\frac{1}{2}},$
which tends to 00 as $n$ goes to infinity. Moreover, $\theta_{n}$ converges pointwise to $1$. Hence, since the pair $(G,H)$ has relative Property $\mathrm{(T_{P})}$, there is $n_{0}\geq 1$ such that
$\sup_{x,y\in H}|\theta_{n_{0}}(x,y)-1|<\frac{1}{2}.$
In other words,
$\sup_{x,y\in H}\|R(x)-R(y)\|^{2}<n_{0}\log(2),$
which implies that, for all $x\in H$,
$\psi(x)$ $=\|R(x)-R(1)\|^{2}+\|S(x)+S(1)\|^{2}$
$<n_{0}\log(2)+\psi(1).$
Since $H$ is infinite and $\psi$ is proper, this is not possible. We conclude that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)>1$. ∎
We can now prove Corollaries 1.3 and 1.4.
Proof of Corollary 1.3.
Let $\Gamma$ be a lattice in $G(\mathbb{K})$. By [3, Corollary 4.8], $\Gamma$ has Property $\mathrm{(TTT)}$. Hence, by Theorem 1.2, $\boldsymbol{\Lambda}_{\mathrm{WH}}(\Gamma)>1$. ∎
Proof of Corollary 1.4.
Let $G=\Gamma\ltimes A$, where $A$ is infinite and abelian, and assume that the pair $(G,A)$ has relative Property $\mathrm{(T)}$. By [9, Proposition 3], $(G,A)$ has relative Property $\mathrm{(TTT)}$. Thus, by Theorem 1.2, $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)>1$. ∎