### v2

# Property $\mathrm{(TTT)}$ and the Cowling–Haagerup constant

Abstract.
We show that Property $\mathrm{(TTT)}$ is an obstruction to weak amenability with Cowling–Haagerup constant $1$. More precisely, if $G$ is a countable group and $H$ is an infinite subgroup of $G$ such that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$, then the weak Haagerup constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ is strictly greater than $1$. We apply this result to some semidirect products and lattices in higher rank algebraic groups.
Key words and phrases:
Property (TTT), weak amenability, weak Haagerup property
2020 Mathematics Subject Classification:
Primary 22D55; Secondary 43A07, 46L07
This work is supported by the FONDECYT project 3230024.
1. Introduction
A countable group $G$ is said to be amenable if there is a sequence of finitely supported, positive definite functions $\varphi_{n}:G\to\mathbb{C}$ such that
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G.$
More generally, we say that $G$ is weakly amenable if there is a sequence of finitely supported functions $\varphi_{n}:G\to\mathbb{C}$ and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G,$
and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C,$
where $B_{2}(G)$ stands for the space of Herz–Schur multipliers of $G$; see Section 2.2. The Cowling–Haagerup constant $\boldsymbol{\Lambda}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds. This is indeed a weakening of amenability because every positive definite function $\varphi$ satisfies $\|\varphi\|_{B_{2}(G)}=\varphi(1)$. Hence, if $G$ is amenable, then $\boldsymbol{\Lambda}(G)=1$. If $G$ is not weakly amenable, we simply set $\boldsymbol{\Lambda}(G)=\infty$. We refer the reader to [17] for a survey on weak amenability.
A strong negation of amenability is given by Property $\mathrm{(T)}$. Let $G$ be a countable group, and let $H$ be a subgroup of $G$. We say that the pair $(G,H)$ has relative Property $\mathrm{(T)}$ if every sequence of positive definite functions on $G$ converging pointwise to $1$ converges uniformly on $H$. If this condition holds for $H=G$, we simply say that $G$ has Property $\mathrm{(T)}$. We refer the reader to [1] for a thorough account on Property $\mathrm{(T)}$ and its different characterisations. It follows from the definitions that, if $H$ is infinite and the pair $(G,H)$ has relative Property $\mathrm{(T)}$, then $G$ cannot be amenable.
For weak amenability, however, the situation is quite different, and Property $\mathrm{(T)}$ is no longer an obstruction. This was first observed by Cowling and Haagerup [2], who showed that the rank 1 Lie groups $\operatorname{Sp}(n,1)$ and $\operatorname{F}_{4,-20}$, as well as their lattices, are weakly amenable. Afterwards, the list of examples was considerably expanded by Ozawa [13], who showed that all hyperbolic groups are weakly amenable. It must be noted that, in a sense that can be made very precise, “almost all” hyperbolic groups have Property $\mathrm{(T)}$; see [18] and [9] for details. In spite of these facts, the following question remains open.
Question 1.1.
Does there exist a pair $(G,H)$ with relative Property $\mathrm{(T)}$ such that $H$ is infinite and $\boldsymbol{\Lambda}(G)=1$?
As a matter of fact, all the known examples of groups with $\boldsymbol{\Lambda}(G)=1$ satisfy the Haagerup property, meaning that there is a sequence of positive definite functions $\varphi_{n}:G\to\mathbb{C}$ such that $\varphi_{n}$ vanishes at infinity for all $n$, and
$\lim_{n\to\infty}\varphi_{n}(s)=1,\quad\forall s\in G.$
The problem of determining whether this implication holds in general is often referred to as Cowling’s conjecture. In order to attack this problem, Knudby [7] introduced the weak Haagerup property as a generalisation of both weak amenability and the Haagerup property. In this case, the approximation of the identity is given by functions vanishing at infinity and which are uniformly bounded in $B_{2}(G)$. We can also associate a constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ to this property, which is defined in similar fashion; see Section 2.2. We always have
$\boldsymbol{\Lambda}_{\mathrm{WH}}(G)\leq\boldsymbol{\Lambda}(G).$
In this paper, we address a variation of Question 1.1 by considering a stronger version of Property $\mathrm{(T)}$. In [14], Ozawa introduced Property $\mathrm{(T_{P})}$ as a non-equivariant version of Property $\mathrm{(T)}$, where positive definite functions are replaced by positive definite kernels, with some additional assumptions; see Section 2.3 for details. An a priori weaker property, called Property $\mathrm{(TTT)}$, which was also introduced in [14], is defined in terms of boundedness of wq-cocycles; see Section 2.3. Very recently, Dumas [3] showed that Property $\mathrm{(T_{P})}$ and Property $\mathrm{(TTT)}$ are equivalent. Using this fact, we prove the following.
Theorem 1.2.
Let $G$ be a countable group, and let $H$ be an infinite subgroup of $G$ such that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$. Then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ belongs to $(1,\infty]$.
Property $\mathrm{(TTT)}$ is strictly stronger than Property $\mathrm{(T)}$, as witnessed by the fact that hyperbolic groups are a-$\mathrm{TTT}$-menable, which is a strong negation of Property $\mathrm{(TTT)}$; see [14]. However, these properties are equivalent for some semidirect products; see [14, Proposition 3]. Using this fact, we obtain the following.
Corollary 1.3.
Let $G$ be a countable group such that $G=\Gamma\ltimes A$, where $A$ is infinite and abelian. If the pair $(G,A)$ has relative Property $\mathrm{(T)}$, then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ belongs to $(1,\infty]$.
In [3, Corollary 4.8], Property $\mathrm{(TTT)}$ was established for lattices in higher rank algebraic groups. Hence, by Theorem 1.2, such lattices satisfy $\boldsymbol{\Lambda}_{\mathrm{WH}}(\Gamma)>1$. It turns out that something much stronger is true. The following result is a consequence of [5], [10] and [11]. Since it has not appeared explicitly in the literature, we include it here together with a proof. For details on algebraic groups, we refer the reader to [12, Chapter 1].
Theorem 1.4.
Let $\mathbb{K}$ be a local field, and let $G$ be a connected, almost $\mathbb{K}$-simple $\mathbb{K}$-group with $\operatorname{rank}_{\mathbb{K}}G\geq 2$. Then $G(\mathbb{K})$ does not have the weak Haagerup property. The same holds for any lattice in it.
This paper is organised as follows. In Section 2, we recall the main definitions and properties studied in this manuscript. In Section 3, we prove Theorem 1.2 and Corollary 1.3. Finally, in Section 4, we discuss algebraic groups, and prove Theorem 1.4.
Acknowledgements
I’m grateful to Mikael de la Salle for his insightful comments and suggestions.
2. Preliminaries
2.1. Schur multipliers
We begin by reviewing some facts about Schur multipliers and positive definite kernels; for more detailed presentations, we refer the reader to [15, Chapter 5] and [1, Appendix C]. Let $X$ be a set. We say that $\theta:X\times X\to\mathbb{C}$ is a Schur multiplier on $X$ if there is a Hilbert space $\mathcal{H}$ and bounded maps $P,Q:X\to\mathcal{H}$ such that
$\theta(x,y)=\langle P(x),Q(y)\rangle,\quad\forall x,y\in X.$ (1)
We denote by $V(X)$ the space of Schur multipliers on $X$, and we endow it with the norm
$\|\theta\|_{V(X)}=\inf\left(\sup_{x\in X}\|P(x)\|\right)\left(\sup_{y\in X}\|Q(y)\|\right),$
where the infimum is taken over all the decompositions as in (1). In the particular case in which (1) can be obtained with $P=Q$, we say that $\theta$ is a positive definite kernel on $X$. In addition, we say that $\theta$ is normalised if $\theta(x,x)=1$ for all $x\in X$. We will make use of the following basic fact; see [1, Theorem C.3.2].
Proposition 2.1.
Let $\xi:X\to\mathcal{H}$ be a map from a set $X$ to a Hilbert space $\mathcal{H}$. For every $t>0$, the map
$(x,y)\in X\times X\ \longmapsto\ e^{-t\|\xi(x)-\xi(y)\|^{2}}$
is a normalised, positive definite kernel on $X$.
2.2. Weak amenability and the weak Haagerup property
Now let $G$ be a locally compact, second countable group. We say that a continuous function $\varphi:G\to\mathbb{C}$ is a Herz–Schur multiplier if the map
$(x,y)\in G\times G\ \longmapsto\ \varphi(y^{-1}x)$
is a Schur multiplier on $G$. We denote by $B_{2}(G)$ the space of Herz–Schur multipliers on $G$, and we endow it with the norm inherited from $V(G)$, for which it becomes a Banach algebra. We say that $G$ is weakly amenable if there is a sequence of compactly supported Herz–Schur multipliers $\varphi_{n}\in B_{2}(G)$ and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\sup_{x\in K}|\varphi_{n}(x)-1|=0,$
for every compact subset $K\subseteq G$, and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C.$
The Cowling–Haagerup constant $\boldsymbol{\Lambda}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds.
More generally, $G$ has the weak Haagerup property if there is a sequence $\varphi_{n}\in B_{2}(G)$ such that $\varphi_{n}$ vanishes at infinity for all $n$, and a constant $C\geq 1$ such that
$\lim_{n\to\infty}\sup_{x\in K}|\varphi_{n}(x)-1|=0,$
for every compact subset $K\subseteq G$, and
$\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{B_{2}(G)}\leq C.$
The weak Haagerup constant $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)$ is the infimum of all $C\geq 1$ such that the condition above holds. It readily follows that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)\leq\boldsymbol{\Lambda}(G)$. We point out that these constants may very well be different. Even more, there are groups $G$ such that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$ and $\boldsymbol{\Lambda}(G)=\infty$. One example of such a group is the wreath product $\mathbb{Z}\wr\mathbb{F}_{2}$, where $\mathbb{F}_{2}$ is the free group on 2 generators; see [8] for details.
The weak Haagerup property was introduced in [7] with the goal of studying the connection between weak amenability and the Haagerup property. The following result will be essential to our purposes; see [7, Theorem 1.2] and [7, Proposition 4.3].
Theorem 2.2 (Knudby).
Let $G$ be a countable discrete group with $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$. Then there exist a function $\psi:G\to[0,\infty)$, a Hilbert space $\mathcal{H}$, and maps $R,S:G\to\mathcal{H}$ such that
$\psi(y^{-1}x)=\|R(x)-R(y)\|^{2}+\|S(x)+S(y)\|^{2},\quad\forall x,y\in G,$ (2)
and $\lim_{x\to\infty}\psi(x)=\infty$.
Observe that the map $S:G\to\mathcal{H}$ given by Theorem 2.2 takes values in the sphere of radius $\frac{1}{2}\sqrt{\psi(1)}$ in $\mathcal{H}$.
2.3. Property $\mathrm{(T_{P})}$ and Property $\mathrm{(TTT)}$
Now we discuss two strengthenings of Property $\mathrm{(T)}$ introduced in [14]. For simplicity, we shall restrict ourselves to discrete groups. Let $G$ be a group and let $\theta:G\times G\to\mathbb{C}$. For every $s\in G$, we will denote by $s\cdot\theta$ the map
$s\cdot\theta(x,y)=\theta(s^{-1}x,s^{-1}y),\quad\forall x,y\in G.$
Now let $H$ be a subgroup of $G$. We say that the pair $(G,H)$ has relative Property $\mathrm{(T_{P})}$ if, for every $\varepsilon>0$, there is a finite subset $K\subset G$ and $\delta>0$ such that, for every normalised positive definite kernel $\theta:G\times G\to\mathbb{C}$ satisfying
$\sup_{s\in G}\|s\cdot\theta-\theta\|_{V(G)}$ $<\delta$ and $\sup_{x^{-1}y\in K}|\theta(x,y)-1|<\delta,$
one has
$\sup_{x,y\in H}|\theta(x,y)-1|<\varepsilon.$
Let $\mathcal{H}$ be a Hilbert space and let $\mathbf{U}(\mathcal{H})$ denote its unitary group. We say that $b:G\to\mathcal{H}$ is a wq-cocycle if there is a map $\pi:G\to\mathbf{U}(\mathcal{H})$ (not necessarily a representation) such that
$\sup_{x,y\in G}\|b(xy)-\pi(x)b(y)-b(x)\|<\infty.$
We say that the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$ if every wq-cocycle on $G$ is bounded on $H$. If $H=G$, we simply say that $G$ has Property $\mathrm{(TTT)}$. The following was proved in [14, Theorem 1] and [3, Theorem 2.7].
Theorem 2.3 (Ozawa, Dumas).
Let $G$ be a countable group, and let $H$ be a subgroup of $G$. Then the pair $(G,H)$ has relative Property $\mathrm{(TTT)}$ if and only if it has relative Property $\mathrm{(T_{P})}$.
3. Proof of the main result
Now we prove Theorem 1.2. The proof is inspired by that of [14, Theorem 1].
Proof of Theorem 1.2.
Let $G$ be a countable group, and let $H$ be an infinite subgroup of $G$ such that $(G,H)$ has relative Property $\mathrm{(TTT)}$. By Theorem 2.3, $(G,H)$ has relative Property $\mathrm{(T_{P})}$. Assume by contradiction that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=1$, and let $R,S:G\to\mathcal{H}$ be the maps given by Theorem 2.2. For every $n\geq 1$, we define $\theta_{n}:G\times G\to(0,1]$ by
$\theta_{n}(x,y)=e^{-\frac{1}{n}\|R(x)-R(y)\|^{2}},\quad\forall x,y\in G.$
By Proposition 2.1, $\theta_{n}$ is a normalised positive definite kernel for every $n\geq 1$. Observe that we can view $\mathcal{H}$ as a real Hilbert space with inner product $\mathrm{Re}(\langle\cdot,\cdot\rangle)$. Now consider the full Fock Hilbert space
$\mathcal{F}=\bigoplus_{n\geq 0}\mathcal{H}^{\otimes n},$
where $\mathcal{H}^{\otimes 0}={\mathbb{R}}$ and $\mathcal{H}^{\otimes(n+1)}=\mathcal{H}\otimes\mathcal{H}^{\otimes n}$. The exponential map $\operatorname{EXP}:\mathcal{H}\to\mathcal{F}$ is given by
$\operatorname{EXP}(\xi)=1\oplus\frac{\xi}{\sqrt{1!}}\oplus\frac{\xi\otimes\xi}{\sqrt{2!}}\oplus\frac{\xi\otimes\xi\otimes\xi}{\sqrt{3!}}\oplus\cdots,\quad\forall\xi\in\mathcal{H}.$
We can define a map from $\mathcal{H}$ to the unit sphere of $\mathcal{F}$ by
$\operatorname{E}(\xi)=e^{-\|\xi\|^{2}}\operatorname{EXP}(\sqrt{2}\xi),\quad\forall\xi\in\mathcal{H}.$
This map satisfies
$\langle\operatorname{E}(\xi),\operatorname{E}(\eta)\rangle=e^{-\|\xi-\eta\|^{2}},\quad\forall\xi,\eta\in\mathcal{H}.$
For each $n\geq 1$, define $P_{n},Q_{n}:G\to\mathcal{F}$ by
$P_{n}(x)$ $=\operatorname{E}\left(\frac{1}{\sqrt{n}}S(x)\right),$ $Q_{n}(x)$ $=\operatorname{E}\left(-\frac{1}{\sqrt{n}}S(x)\right),$
for all $x\in G$. Recall the identity (2) and observe that, for all $s,x,y\in G$, $n\geq 1$,
$\theta_{n}(x,y)-s\cdot\theta_{n}(x,y)$ $=e^{-\frac{1}{n}\|R(x)-R(y)\|^{2}}-e^{-\frac{1}{n}\|R(s^{-1}x)-R(s^{-1}y)\|^{2}}$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\left(e^{\frac{1}{n}\|S(x)+S(y)\|^{2}}-e^{\frac{1}{n}\|S(s^{-1}x)+S(s^{-1}y)\|^{2}}\right)$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\left(\langle P_{n}(x),Q_{n}(y)\rangle-\langle P_{n}(s^{-1}x),Q_{n}(s^{-1}y)\rangle\right)$
$=e^{-\frac{1}{n}\psi(y^{-1}x)}\langle P_{n}(x)-P_{n}(s^{-1}x),Q_{n}(y)\rangle$
$\quad+e^{-\frac{1}{n}\psi(y^{-1}x)}\langle P_{n}(s^{-1}x),Q_{n}(y)-Q_{n}(s^{-1}y)\rangle.$
Furthermore,
$\|P_{n}(x)-P_{n}(s^{-1}x)\|^{2}$ $=2-2\langle P_{n}(x),P_{n}(s^{-1}x)\rangle$
$=2-2e^{-\frac{1}{n}\|S(x)-S(s^{-1}x)\|^{2}}$
$\leq 2-2e^{-\frac{1}{n}\psi(1)}.$
Similarly, we get
$\|Q_{n}(y)-Q_{n}(s^{-1}y)\|^{2}\leq 2-2e^{-\frac{1}{n}\psi(1)}.$
On the other hand, by [7, Proposition 4.3],
$\left\|(x,y)\mapsto e^{-\frac{1}{n}\psi(y^{-1}x)}\right\|_{V(G)}\leq 1.$
All this shows that
$\sup_{s\in G}\|s\cdot\theta_{n}-\theta_{n}\|_{V(G)}\leq 2\sqrt{2}\left(1-e^{-\frac{1}{n}\psi(1)}\right)^{\frac{1}{2}},$
which tends to 00 as $n$ goes to infinity. On the other hand, for every finite subset $K\subset G$, we can define
$M_{K}=\sup_{x\in K}\psi(x),$
which gives us
$\sup_{x^{-1}y\in K}\|R(x)-R(y)\|^{2}\leq M_{K}.$
Equivalently,
$\sup_{x^{-1}y\in K}|\theta_{n}(x,y)-1|\leq 1-e^{-\frac{M_{K}}{n}},$
which again tends to 00 as $n$ goes to infinity. Hence, since the pair $(G,H)$ has relative Property $\mathrm{(T_{P})}$, there is $n_{0}\geq 1$ such that
$\sup_{x,y\in H}|\theta_{n_{0}}(x,y)-1|<\frac{1}{2}.$
In other words,
$\sup_{x,y\in H}\|R(x)-R(y)\|^{2}<n_{0}\log(2),$
which implies that, for all $x\in H$,
$\psi(x)$ $=\|R(x)-R(1)\|^{2}+\|S(x)+S(1)\|^{2}$
$<n_{0}\log(2)+\psi(1).$
Since $H$ is infinite and $\psi$ is proper, this is not possible. We conclude that $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)>1$. ∎
We can now prove Corollary 1.3.
Proof of Corollary 1.3.
Let $G=\Gamma\ltimes A$, where $A$ is infinite and abelian, and assume that the pair $(G,A)$ has relative Property $\mathrm{(T)}$. By [14, Proposition 3], $(G,A)$ has relative Property $\mathrm{(TTT)}$. Thus, by Theorem 1.2, $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)>1$. ∎
4. Higher rank algebraic groups
Now we discuss algebraic groups and the proof of Theorem 1.4. We quickly recall the basic definitions that we will need; for a detailed treatment, we refer the reader to [12, Chapter 1].
Let $\mathbb{K}$ be a local field, and let $G$ be a $\mathbb{K}$-group. We say that $G$ is almost $\mathbb{K}$-simple if every $\mathbb{K}$-closed normal subgroup of $G$ is finite. Let $\operatorname{rank}_{\mathbb{K}}G$ denote the $\mathbb{K}$-rank of $G$, i.e. the common dimension of maximal $\mathbb{K}$-split tori in $G$. We also denote by $G(\mathbb{K})$ the set of $\mathbb{K}$-rational points of $G$, which is a locally compact group. The following was proved in [3, Corollary 4.8].
Theorem 4.1 (Dumas).
Let $\mathbb{K}$ be a local field, and let $G$ be a connected, almost $\mathbb{K}$-simple $\mathbb{K}$-group with $\operatorname{rank}_{\mathbb{K}}G\geq 2$. If $\Gamma$ is a lattice in $G(\mathbb{K})$, then $\Gamma$ has Property $\mathrm{(TTT)}$.
As a consequence, we obtain the following.
Corollary 4.2.
Let $\Gamma$ be as in Theorem 4.1, then $\boldsymbol{\Lambda}_{\mathrm{WH}}(\Gamma)>1$.
Proof.
By Theorem 4.1, $\Gamma$ has Property $\mathrm{(TTT)}$. Hence, by Theorem 1.2, $\boldsymbol{\Lambda}_{\mathrm{WH}}(\Gamma)>1$. ∎
As mentioned in the introduction, we actually have $\boldsymbol{\Lambda}_{\mathrm{WH}}(\Gamma)=\infty$ for such lattices. This is a consequence of a series of different results that we record here for the reader’s convenience.
In [4], Haagerup showed that higher rank simple Lie groups with finite centre are not weakly amenable. The proof focuses on the groups $\operatorname{SL}_{3}({\mathbb{R}})$ and $\operatorname{Sp}_{4}({\mathbb{R}})$, and the general result follows from these two particular cases. This strategy has been adapted to other similar contexts; see [5, 6, 11, 16]. The proof of Theorem 1.4 follows the same ideas.
Let $\mathbb{K}$ be a field. The special linear group $\operatorname{SL}_{3}(\mathbb{K})$ is the group of $3\times 3$ matrices with coefficients in $\mathbb{K}$ and determinant 1. The symplectic group $\operatorname{Sp}_{4}(\mathbb{K})$ is given by all matrices $A\in\operatorname{GL}_{4}(\mathbb{K})$ such that $A^{t}JA=J$, where
$J=\left(\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&-1&0&0\\ \end{array}\right).$
The following was proved in [10, Proposition 4.1] and [10, Proposition 5.1].
Theorem 4.3 (Lafforgue–de la Salle).
Let $\mathbb{K}$ be either ${\mathbb{R}}$ or a nonarchimedian local field. Then $\boldsymbol{\Lambda}_{\mathrm{WH}}(\operatorname{SL}_{3}(\mathbb{K}))=\infty$.
Remark 4.4.
The results in [10] are expressed in a different (more general) language. They involve the spaces $MS^{p}(L^{2}(G))$ of multipliers of Schatten classes. Here we are only interested in the case $p=\infty$, for which $\|\check{\varphi}\|_{MS^{\infty}(L^{2}(G))}=\|\varphi\|_{B_{2}(G)}$; see [10, §1] for details.
In the real case, this result was extended in [6] to the symplectic group $\operatorname{Sp}_{4}({\mathbb{R}})$ (where it is called $\operatorname{Sp}(2,{\mathbb{R}})$). Using this, the following was obtained in [5, Theorem B].
Theorem 4.5 (Haagerup–Knudby).
Let $G$ be a connected simple Lie group with $\operatorname{rank}_{\mathbb{R}}G\geq 2$. Then $\boldsymbol{\Lambda}_{\mathrm{WH}}(G)=\infty$.
In the nonarchimedian case, a similar result for $\operatorname{Sp}_{4}(\mathbb{K})$ was obtained in [11, Theorem 2.4]. Again, the language is quite different, but a particular case of that result can be stated as follows.
Theorem 4.6 (Liao).
Let $\mathbb{K}$ be a nonarchimedian local field. Then $\boldsymbol{\Lambda}_{\mathrm{WH}}(\operatorname{Sp}_{4}(\mathbb{K}))=\infty$.
All these results together allow us to obtain the same conclusion for more general higher rank algebraic groups.
Proof of Theorem 1.4.
Let $G$ be a connected, almost $\mathbb{K}$-simple $\mathbb{K}$-group with $\operatorname{rank}_{\mathbb{K}}G\geq 2$. By [12, Proposition 1.6.2] and [12, Proposition 2.3.4], $G(\mathbb{K})$ has a closed subgroup $F$ which is isomorphic to a quotient of either $\operatorname{SL}_{3}(\mathbb{K})$ or $\operatorname{Sp}_{4}(\mathbb{K})$ by a finite subgroup. Hence, by [8, Theorem A.(1)],
$\boldsymbol{\Lambda}_{\mathrm{WH}}(G(\mathbb{K}))\geq\boldsymbol{\Lambda}_{\mathrm{WH}}(F).$
Moreover, by [8, Theorem A.(2)], together with Theorems 4.3, 4.5 and 4.6, we get
$\boldsymbol{\Lambda}_{\mathrm{WH}}(F)=\infty.$
This shows that $G(\mathbb{K})$ does not have the weak Haagerup property. By [8, Theorem A.(6)], the same holds for any lattice in $G(\mathbb{K})$. ∎